What is the equation of a line with a slope of 6/7 abd passes through (4,-1)

Answer by nerdybill(7384)

What is the equation of a line with a slope of 6/7 and passes through (4,-1)
.
Since you are given "slope" and a "point", you can plug it into the "point-slope" form of a line:
y - y1 = m(x - x1)
y - (-1) = (6/7)(x - 4)
y + 1 = (6/7)(x - 4)
y + 1 = (6/7)x - (6/7)4
y + 1 = (6/7)x - 24/7
y = (6/7)x - 24/7 - 1
y = (6/7)x - 24/7 - 7/7
y = (6/7)x - 31/7 (answer in "slope-intercept" form)